Define the units of specific heat in terms of calories, kilocalories, and joules
Answer #1

We know that,

Specific heat is amount of heat required to raise the temperature per unit mass.

lt means lt is heat required in Joule to raise the temperature by 1 Kelvin of 1 kg mass.

So, unit of specific heat = J / kg.K or J/g.oC

Or

lt is heat required in Calories to raise the temperature by 1 oC of 1 g mass.

So, unit = Cal / g.oC

Or

lt is heat required in Kilojoules to raise the temperature by 1 Kelvin of 1 kg mass.

Unit = KJ / kg.K